1.A company has purchased an asset for $165,461. If they require a return of 11.38%, how much must they sell the asset for in 8 years?(Round to 2 decimal places.)

2.Suppose that 2 years ago you bought an old record player at a yard sale for $4. You saw today on E-bay that the same record player is selling for $72. If you were to sell the record play at that price today, what would be the implied return percentage? (Convert to a percent. Round to 2 decimal places.

3.An investment will pay $17,747 in 16 years. If the interest rate is 13.8%, how much is it worth today? (Round to 2 decimal places.)

4.If you have $462 today and you think you can earn an interest rate of 6% per year, how many years until you have $844?

5.An investment will pay you $3,096 in 1 years if you pay $1,980 today. What is the implied rate of return? (Convert to a decimale. Round to 2 decimal places.)

6.In 1998, the average price of a gallon of gas was $1.09. Today, the average price of a gallon of gas is $2.84. At what annual rate has a gallon of gas increased over the last 20 years? (Answer as a percent. Enter only numbers and decimals in your response. Round to 2 decimal places.)

7.Your goal is to have $76,023. If you can earn 11.9% per year and you invest $18,064 today, how many years until you reach your goal?

8.You are buying an investment property for $383,240 today. Through research, you have determined that the value of the property is likely to grow at a rate of 7.17% per year, on average. How long until the value of the property grows to $521,360?

9.Calculate the value of the cash flows at time 6. The interest rate is 11.99%.

0 1 2
$1,135 $2,375 $3,065

10.You are looking at an investment that will pay you $25,919 in year 2, $42,812 in year 4 and $45,841 in year 6. If your required return is 8.34%, what is the most you should pay for the investment? (In other words, how much is the project worth today?)
Answer #1
1 Assets cost = $165,461
Rate of return (r)= 11.38%
No. of years (n)= 8
Future value = Present value*(1+r)^n
Future value = 165461*(1+11.38%)^8
Future value = 165461*2.3684144
Future value = 391,880.23
2 Purchase value = $4
Future value = $72
No. of years (n) = 2
Implied rate of return = (futureprice/purchase price)^(1/n)-1
Implied rate of return = ((72/4)^1/2)-1
Implied rate of return = 8%
3 Future value = $17747
No. of years (n) = 16
Rate = 13.8%
Present value = 17747*Present value factor at 13.8% for 16 year
Present value = 17747*Present value factor at 13.8% for 16 year
Present value = 17747*0.1554
Present value = 17747*0.1554
Present value = 2757.88
4 Purchase value = $462
Future value = $844
No. of years (n)
Rate of interest = 6%
Future value = Present value*(1+r)^n
844=462*(1+6%)^n
844/462 =(1+6%)^n
1.82684 at 6% from Present value annuity factor
No. of years = 1.993
No. of years = 2